# Outward Sodium Current, Fine-structure Constant, and Ferroelectric Hysteresis Regimes in the Giant Squid Axon Propagating Action Potential: a Phase Space Approach

### v8

@@ III Analysis - Part 2
where $T$ is the temperature in Kelvin degrees, $\kappa_{B}$ is the Boltzmann constant. $\epsilon_{X}$ and $\kappa_{X}$ are respectively the activation energy and the Boltzmann time rate. See Fig. 13 for graphs of Eqs. (23) for the ions traversing sodium and potassium channels and for the propagation constant $k$. All fits are linear for up to about $20^{o}$.
@@
Figure 16: Rising edge mAvrami fits of fraction of open $Mo/M$ sodium M-channels and fits of polarization channels fractions $Pio/Pi$. Inception and AP peak polarization flips are displayed. Parameters $\alpha_{M}$ and $\alpha_{Pi}$ from Eq.(20) are seeded with the value of the fine-structure constant $\alpha$ = 0.007297352 and parameters $\theta_{M}$, $\theta_{Pi}$ are seeded with the value 3.78. Note: $g_{P1}=g_{P2}$, $V_{P1}=V_{P2}$, $g_{P3}=g_{P4}$ and $V_{P3}=V_{P4}$. (a) Both, the inception and AP peak polarization segments consist of two concatenated portions with different time rates. Inception polarization flip interpolation: + . Note, that polarization time rates $\mu_{P2,3,4}$ are close multiples of sodium’s M-channel time rate $\mu_{M}$. See SM Fig. 25 where all four $\mu_{Pi}$ are close multiples of $\mu_{M}$. (b), (c), (d) Note: $\ln\alpha$ = $\ln(0.007297352...)$ = $-4.920243...$.
Figure 17: Recovery mAvrami fits of fraction of open sodium H-channels $Ho/H$ and potassium N-channels $No/N$. Parameters $\alpha_{H}$ and $\alpha_{N}$ from Eq.(20) are seeded with the value of the fine-structure constant $\alpha$ = 0.007297352 and parameters $\theta_{H}$, $\theta_{N}$ are adjusted by fitting. (a),(c) Plots of experimental fractions $Ho(t)/H$ and $No(t)/N$ vs. time[mSec] are fitted with mAvrami Eq.(25b) and Eq.(25c) respectively. (b), (d) Plots of experimental $\ln[-\ln(1-Ho(t)/H)]$ vs. $\ln[\mu_{H}(to_{H}-t)]$ and $\ln[-\ln(1-No(t)/N)]$ vs. $\ln[\mu_{N}(t-to_{N})]$ are fitted with linear functions. Note: $\ln\alpha$ = $\ln(0.007297352...)$ = $-4.920243...$.
@@
where $to_{M}$ is the inception time, the time when sodium channels start opening, and $A_{M}$ and $\theta_{M}$ are Avrami parameters. The Avrami equation is best known for describing isothermal phase changing in solids and crystallization processes. The particular value of the exponent $\theta=4$ is said to have contributions from three dimensions of growth and one representing a constant nucleation rate. Avrami parameter $A_{X}$ is typically very temperature dependent. In particular, for incoming sodium current, $A_{M}$ varies exponentially from about 60 at $1^{o}$ to 15000 at $25^{o}$ (See SM Fig. 20 for the table of parameters.)
@@
where $to_{M}$ and $to_{N}$ are the temperature dependent times when sodium M and potassium N channels start to open, $tc_{H}$ is the time when sodium H channels close, $to_{P1,3,4}$ are the times when polarization segments $P{1,3,4}$ channels open and $tc_{P2}$ is the time when the segment $P2$ channels close. We did not attempt the detailed fitting of Ko/K. The very small ionic current at the foot of action potential is a dynamical equilibrium of potassium, chloride and sodium currents.
@@
Incidentally, there is a linear relationship between 1/$\mu_{M}$, 1/$\mu_{P4}$ and $to_{M}$, $t_{oP4}$ respectively, see SM Fig. 27. Knowing the time rate $\mu_{M}$ (or conductance $g_{M}$), the relationship predicts the total time $|to_{M}|$ to open all M-channels or vice versa.
@@
The current $J_{M}(V)$ is about three times larger at 19.8 oC than at 4.5 oC since $g_{M}$ is about three times larger at 19.8 oC than at 4.5 oC and the factor $[Mo(V)/M](V-V_{M})$ is moderately temperature dependent. And, M channels open completely in a time interval about three times shorter at 19.8 oC than at 4.5 oC. Therefore, the total number of incoming sodium ions, at 4.5 oC and up to 19.8 oC, is about the same at all temperatures.
[…]
The ionic polarization current has two stages in the post inception polarization segment, both with same maximum conductance $g_{K}$ <<much-less-than$<$ < $g_{P1}=g_{P2}$ <<much-less-than$<$ < $g_{M}$ and with same reversal potential $V_{P1}=V_{P2}$ <<much-less-than$<$ < $V_{M}$. The polarization segment with the time rate $\mu_{P1}$ starts with the channel symmetry polarization flip from H-symmetry to M-symmetry. The second segment with the time rate $\mu_{P2}$ follows shutting down the inception ionic polarization current. Among 11 sweeps at different temperatures, only Sweep170 appears to close the inception ionic polarization current via two different time rates. We have fitted only the first one (See SM Fig. 25. For a typical sweep see Fig. 16.
@@
where $V_{p}$ is the value of the potential at the peak of the action potential, $Vo_{M}$ and $Vo_{N}$ are the respective potentials at which sodium M-channels and potassium N-channels start to open, $Vc_{H}$ is the potential at which sodium H-channels close, and $Vo_{P3,4}$ is the potential at which the polarization $P{3,4}$ channels start opening. Fits for fractions $Mo(V)/M$ are seeded with the value $Vo_{M}$ corresponding to $to_{M}$. The average value of the $\delta_{M}$ exponent for ten sweeps from $1^{o}$ to $25.2^{o}$ is $\delta_{M}$ = 3.07. Fits for fractions Ho(V)/H and No(V)/N have been done with values $Vo_{H}$ and $Vo_{N}$ as free parameters. Note that the fractions of open channels $Mo(V)/M$, $Ho(V)/H$ and $No(V)/N$ at any given time can be interpreted as the fraction of time that the corresponding single M-channel, H-channel or N-channel is stochastically open. The corresponding phase space ionic currents fits for Sweep 170 are plotted in Fig. 21 and Fig. 8. AP peak polarization segments, with steep tangents in the proximity of the peak and with more than one time rate, are less amenable to phase space fitting.
@@ IV Discussion and Conclusion
Our phenomenological fittings of experimental currents associated with propagating action potentials—using pertinent physical parameters—reveal the universal role of the fine-structure constant in ion channel conduction. These fittings are notably accurate across the full range of biological temperatures, suggesting that the resulting conclusions and predictions must, in some fashion, be true.
In 1900, Max Planck postulated—without any theoretical foundation—that electromagnetic energy could be emitted only in discrete quantities. He introduced this idea solely to match experimental data for the full spectrum of blackbody radiation. And yet, because the postulate fit the data so well, logic compels us to consider it, in some essential way, true.
@@
We speculate that the full theoretical description consistent with our fittings will take into account the interaction of traversing ions with the two dimensional atomic surface of the ionic pores, a realm of physics yet to be explored. The precedent has already been established in experiments with two dimensional graphene films with photons where the fine-structure constant can be detected ”practically with the naked eye” [39].
@@
1. Phase Space cable equation. We have derived the Phase Space cable equation for the propagating action potential and as discussed in Appendix A, explicitly defined ionic currents as residuals of the membrane current after subtracting the capacitive current and clarified the dual role of chemical and electrical potentials, enhancing both physical understanding and theoretical consistency. We showed that the three currents of the cable equation and Ohm’s current in the axoplasm are proportional to $\Phi(V)=dV/dt$. We have also shown that the total ionic current crosses the zero current axis at two points when $\Phi(V)/dV=k$ where electrostatic driving force annuls the ionic density gradient. Plots of the three currents exhibit the functional complexity of the membrane current in perpetuating the propagation and keeping the shape of the action potential. In particular, the plot of currents in the action potential frame exhibits the role of the Ohm’s current in accumulating positive charges near the mouths of sodium channels increasing the electrostatic gradient that overwhelms sodium’s concentration gradient resulting in theboutward flow of sodium ions through open sodium channels during recovery.
@@
2. Currents at the foot of the action potential. While the present work deals only with the steady propagation of the action potential, the phase space cable equation provides, without any additional assumption, an insight into the excitation itself. The propagation of the advancing action potential foot is generated by the axoplasmic current $J_{m}$, or more precisely by the product of Ohm’s law current and its derivative increasing the positive charge density at the foot of the action potential and increasing the membrane’s capacitance . This process is equivalent to injecting a charge into the axoplasm to elicit the action potential. The increasing potential at the foot triggers the inception of sodium activation currents and thus perpetuates its propagation. The eventual assumption that sodium activation current $J_{M}$ is an mAvrami current leads to a conclusion that its inception is triggered by a continuous phase change involving the changing of sodium channels symmetry. This fact, for example, suggests that the release of acetylcholine by the action potential at the end of the nerve initiates the signal at the connecting nerve by triggering a continuous phase change. While the present work focuses on the steady-state propagation of the action potential, the phase-space cable equation offers, without any further assumptions, a window into the initiation of excitation itself. The advancing foot of the action potential is driven by the axoplasmic current $J_{m}$—more specifically, by the product of the Ohmic current and its spatial derivative, which increases the positive charge density at the foot and, in turn, the membrane’s capacitive potential. This mechanism is functionally equivalent to injecting charge into the axoplasm to initiate the action potential. The resulting rise in potential at the foot region triggers the onset of sodium activation currents, thereby sustaining the forward propagation of the signal.
If one further assumes that the sodium activation current $J_{M}$ is described by a modified Avrami (mAvrami) process, it follows that its onset corresponds to a continuous phase transition, reflecting a change in the symmetry of sodium channels. This insight suggests, for example, that acetylcholine release at a synapse—initiated by the arrival of an action potential—may trigger excitation in the adjoining neuron by inducing a continuous phase change analogous to that seen in channel activation.
@@
In our model the perturbation that starts the inception of incoming ionic sodium current $J_{M}$ is a continuous phase change with a polarization flip and the inception of the mAvrami activation current $J_{M}$, large capacitive and membrane polarization currents and a small ionic polarization current which depolarizes the membrane after crossing the zero current axis furthering the opening of sodium M-channels and thus increasing the $J_{M}$ incoming sodium current. We have fitted the ionic polarization current leading to the AP peak with some of sodium current $J_{M}$ signature properties: same Avrami exponent 3.78, time rates that are multiples of sodium current time rate $\mu_{M}$, conductance $g_{P3}$ = $g_{P4}$ $\approx g_{M}$ if one excludes the narrow region of the polarization flip, the activation energy $\epsilon_{M}$ slightly smaller than $\epsilon_{P4}$, the reversal potential $V_{P3}$ = $V_{P4}$ $<$ $V_{p}$ but somewhat close to $V_{M}$. The AP peak continuous phase change takes the sodium M-channel symmetry into sodium H-channel symmetry. It follows that the inception AP polarization flip must take sodium H-channel symmetry into sodium M-channel symmetry triggering the action potential. The fitted fraction of open inception ionic polarization channels has: same Avrami exponent 3.78 of M-channel fraction of open channels $Mo(t)/M$, and time rates rates $\mu_{P1}$ and $\mu_{P2}$ at most temperatures are approximate multiples of sodium’s current time rate $\mu_{M}$. However, the conductance $g_{P1,2}$ is much smaller than $g_{M}$. The effective reversal potential $V_{P1}$ = $V_{P2}$ is slightly higher than the inception potential of about 40 mV at $4.5^{o}C$, increasing as the temperature increases. Also, note, that this polarization current changes from outward to incoming as the membrane polarization current crosses the capacitive polarization current. The chloride reversal potential in the axon at rest is reported to be about 15 mV. One should also probably expect a small contribution from polarization channels below the inception potential while in dynamical equilibrium with other players in the AP foot: potassium and sodium permeability, chloride channels current and sodium-potassium pump. However, the present analysis doesn’t provide further insight into the details of the $J_{K}$ current components.
@@
Rosenthal-Bezanilla [43] data covers a range of temperatures from $1^{o}$ to $30^{o}$ yielding the exponential temperature dependence for parameters of interest in our phenomenological theory. Passage of ions through a membrane channel involves interactions with the surface of the pore at microscopic distances. Plots of ionic time rates as Boltzmann’s kinetic rates yield temperature independent Arrhenius activation energies for the passage of ions through the membrane (See Fig. 13). These energies are of the same order as the average of 0.65 eV rate-limiting metabolic biochemical reactions [15]. Traditional Arrhenius processes rely on thermal activation, where temperature plays a direct role in overcoming an energy barrier. Our observed temperature independent Arrhenius rates suggest that thermal fluctuations are not the driving factor in the passage of ions through channels. If the activation energy (typically defined by an Arrhenius relationship) appears constant and unchanging with temperature, it might suggest an alternative mechanism, such as: $Quantum$ $tunneling$: This would allow ions to ”tunnel” through energy barriers without requiring thermal energy. It can be argued that the tunneling rates are equal to ionic time rates $\mu_{X}$. $Fixed$ $field$ or $electromagnetic$ $activation$: In the presence of a strong electric field, ion movement might be governed by field strength rather than thermal fluctuations. In these cases, it’s conceivable that $\alpha$, as suggested by our fittings, as a measure of the electromagnetic interaction strength, could be more relevant than thermal energy in setting the ”activation” characteristics for the process.
@@
3. The ions traverse ionic pores interacting with the surface of the pore at microscopic distances. Our fittings of the fraction of open channels with mAvrami equations yield the phenomenological role of the fine-structure constant, the Arrhenius activation energies that are independent of temperature and the role of the temperature dependent ionic time rates. These facts point to the possibility of an explicit quantum theory justifying the empirical mAvrami equation.
Presented results will provide a framework for further experimental analysis of excitability role by sodium channels lattice and its thermodynamic phase changing behavior, the role of quantum mechanics mediating the flow of ions across ion channels and the plasticity of neuronal stable states essential for memory storage and retrieval mediated by continuous phase changes.

### v9

@@ III Analysis - Part 2
where $T$ is the temperature in Kelvin degrees, $\kappa_{B}$ is the Boltzmann constant. $\epsilon_{X}$ and $\kappa_{X}$ are respectively the activation energy and the Boltzmann time rate. See Fig. 13 for graphs of Eqs. (23) for the ions traversing sodium and potassium channels and for the propagation constant $k$. All fits are linear for up to about $20^{o}$C.
@@
[…]
@@
where $to_{M}$ is the inception time, the time when sodium channels start opening, and $A_{M}$ and $\theta_{M}$ are Avrami parameters. The Avrami equation is best known for describing isothermal phase changing in solids and crystallization processes. The particular value of the exponent $\theta=4$ is said to have contributions from three dimensions of growth and one representing a constant nucleation rate. Avrami parameter $A_{X}$ is typically very temperature dependent. In particular, for incoming sodium current, $A_{M}$ varies exponentially from about 60 at $1^{o}$C to 15000 at $25^{o}$C (See SM Fig. 20 for the table of parameters.)
@@
where $to_{M}$ and $to_{N}$ are the temperature dependent times when sodium M and potassium N channels start to open, $tc_{H}$ is the time when sodium H channels close, $to_{P1,3,4}$ are the times when polarization segments P1,3,4P{1,3,4} channels open and $tc_{P2}$ is the time when the segment $P2$ channels close. We did not attempt the detailed fitting of Ko/K. The very small ionic current at the foot of action potential is a dynamical equilibrium of mainly potassium, and some chloride and sodium currents.
@@
Incidentally, there is a linear relationship between $\mu_{M}, 1$$\mu_{P4}$ and $to_{M}$, $t_{oP4}$ respectively, see SM Fig. 27. Knowing the time rate $\mu_{M}$ (or conductance $g_{M}$), the relationship predicts the total time $|to_{M}|$ to open all M-channels or vice versa.
@@
The current $J_{M}(V)$ is about three times larger at 19.8 oC than at 4.5 oC since $g_{M}$ is about three times larger at 19.8 oC than at 4.5 oC and the factor $Mo(V)/M](V-V_{M})$ is moderately temperature dependent. And, M channels open completely in a time interval about three times shorter at 19.8 oC than at 4.5 oC. Therefore, the total number of incoming sodium ions, at 4.5 oC and up to 19.8 oC, is about the same at all temperatures.
Fig. 16, and SM Figures 25, 26 display, as functions of time, the mAvrami fits for Mo/M and $Pio/Pi$ for $i=1,3,4$ and $P2o/P2$, the fraction of open sodium M-channels and fractions of completed polarization processes at 12.5 oC, 4.5 oC and 14.6 oC respectively. These figures also display the plots of ln[-ln(1-Mo/M)] vs. ln[$\mu_{M}(t-to_{M})]$ and ln[-ln(1-PiPio/PiPi)] vs. ln[$\mu_{Pi}(t-to_{Pi})]$ for $i=1,3,4$ and ln[-ln(1-$P2o/P2$)] vs. ln[$\mu_{P2}(tc_{P2}-t)]$. Note that the fraction of open channels $Mo/M$ at a particular time or corresponding potential is the same as the fraction of time that a single ionic sodium M-channel is open.
The ionic polarization current has two stages in the post inception polarization segment, both with same maximum conductance $g_{K}$ <<<< $g_{P1}=g_{P2}$ <<<< $g_{M}$ and with same reversal potential $V_{P1}=V_{P2}$ <<<< $V_{M}$. The polarization segment with the time rate $\mu_{P1}$ starts with the channel symmetry polarization flip from H-symmetry to M-symmetry. The second segment with the time rate $\mu_{P2}$ follows shutting down the inception ionic polarization current. Among 11 sweeps at different temperatures, only Sweep170 appears to close the inception ionic polarization current via two different time rates. We have fitted only the first one (See SM Fig. 25. For a typical sweep see Fig. 16.
@@
where $V_{p}$ is the value of the potential at the peak of the action potential, $Vo_{M}$ and $Vo_{N}$ are the respective potentials at which sodium M-channels and potassium N-channels start to open, $Vc_{H}$ is the potential at which sodium H-channels close, and $Vo_{P3,4}$ is the potential at which the polarization P3,4P{3,4} channels start opening. Fits for fractions $Mo(V)/M$ are seeded with the value $Vo_{M}$ corresponding to $to_{M}$. The average value of the $\delta_{M}$ exponent for ten sweeps from $1^{o}$C to $25.2^{o}$C is $\delta_{M}$ = 3.07. Fits for fractions Ho(V)/H and No(V)/N have been done with values $Vo_{H}$ and $Vo_{N}$ as free parameters. Note that the fractions of open channels $Mo(V)/M$, $Ho(V)/H$ and $No(V)/N$ at any given time can be interpreted as the fraction of time that the corresponding single M-channel, H-channel or N-channel is stochastically open. The corresponding phase space ionic currents fits for Sweep 170 are plotted in Fig. 21 and Fig. 8. AP peak polarization segments, with steep tangents in the proximity of the peak and with more than one time rate, are less amenable to phase space fitting.
@@ IV Discussion and Conclusion
[…]
Our phenomenological fittings revealing a number of results and predictions bear similarity with Max Planck’s postulate of ”quanta”. In 1900, Max Planck postulated—without any theoretical foundation—that electromagnetic energy could be emitted only in discrete quantities. He introduced this idea solely to match experimental data for the full spectrum of blackbody radiation. And yet, because the postulate fit the data so well, logic compels us to consider it, in some essential way, true.
@@
We speculate that the full theoretical description consistent with our fittings will take into account the interaction of traversing ions with the two dimensional atomic surface of the ionic pores, including tunnelling, a realm of physics yet to be explored. The precedent has already been established in experiments with two dimensional graphene films with photons where the fine-structure constant can be detected ”practically with the naked eye” [39].
In what follows we review and summarize our findings along the action potential trajectory in time:
@@
1. Phase Space cable equation. We have derived the Phase Space cable equation for the propagating action potential and as discussed in Appendix A, explicitly defined ionic currents as residuals of the membrane current after subtracting the capacitive current and clarified the dual role of chemical and electrical potentials, enhancing both physical understanding and theoretical consistency. We showed that the three currents of the cable equation and Ohm’s current in the axoplasm are proportional to $\Phi(V)=dV/dt$. We have also shown that the total ionic current crosses the zero current axis at two points when $d\Phi(V)/dV=k$ where electrostatic driving force annuls the ionic density gradient. Plots of the three currents exhibit the functional complexity of the membrane current in perpetuating the propagation and keeping the shape of the action potential. In particular, the plot of currents in the action potential frame exhibits the role of the Ohm’s current in accumulating positive charges near the mouths of sodium channels increasing the electrostatic gradient that overwhelms sodium’s concentration gradient resulting in the outward flow of sodium ions through open sodium channels during recovery.
@@
2. Currents at the foot of the action potential. While the present work deals only with the steady propagation of the action potential, the phase space cable equation provides, without any additional assumption, an insight into the excitation itself. The propagation of the advancing action potential foot is generated by the axoplasmic current $J_{m}$, or more precisely by the product of Ohm’s law current and its derivative increasing the positive charge density at the foot of the action potential and increasing the capacitive current linearly with the increasing potential. This process is equivalent to injecting a charge into the axoplasm to elicit the action potential. The increasing potential at the foot triggers the inception of sodium activation currents and thus perpetuates its propagation. The eventual assumption that sodium activation current $J_{M}$ is an mAvrami current leads to a conclusion that its inception is triggered by a continuous phase change involving the changing of sodium channels symmetry. This fact, for example, suggests that the release of acetylcholine by the action potential at the end of the nerve initiates the signal at the connecting nerve by triggering a continuous phase change. While the present work focuses on the steady-state propagation of the action potential, the phase-space cable equation offers, without any further assumptions, a window into the initiation of excitation itself. The advancing foot of the action potential is driven by the axoplasmic current $J_{m}$—more specifically, by the product of the Ohmic current and its spatial derivative, which increases the positive charge density at the foot and, in turn, the membrane’s capacitive potential. This mechanism is functionally equivalent to injecting charge into the axoplasm to initiate the action potential. The resulting rise in potential at the foot region triggers the onset of sodium activation currents, thereby sustaining the forward propagation of the signal.
@@
In our model the perturbation that starts the inception of incoming ionic sodium current $J_{M}$ is a continuous phase change with a polarization flip and the inception of the mAvrami activation current $J_{M}$, large capacitive and membrane polarization currents and a small ionic polarization current which depolarizes the membrane after crossing the zero current axis furthering the opening of sodium M-channels and thus increasing the $J_{M}$ incoming sodium current. We have fitted the ionic polarization current leading to the AP peak with some of sodium current $J_{M}$ signature properties: same Avrami exponent 3.78, time rates that are multiples of sodium current time rate $\mu_{M}$, conductance $g_{P3}$ = $g_{P4}$ $\approx g_{M}$ if one excludes the narrow region of the polarization flip, the activation energy $\epsilon_{M}$ slightly smaller than $\epsilon_{P4}$, the reversal potential $V_{P3}$ = $V_{P4}$ << $V_{p}$ but somewhat close to $V_{M}$. The AP peak continuous phase change takes the sodium M-channel symmetry into sodium H-channel symmetry. It follows that the inception AP polarization flip must take sodium H-channel symmetry into sodium M-channel symmetry triggering the action potential. The fitted fraction of open inception ionic polarization channels has: same Avrami exponent 3.78 of M-channel fraction of open channels $Mo(t)/M$, and time rates rates $\mu_{P1}$ and $\mu_{P2}$ at most temperatures are approximate multiples of sodium’s current time rate $\mu_{M}$. However, the conductance $g_{P1,2}$ is much smaller than $g_{M}$. The effective reversal potential $V_{P1}$ = $V_{P2}$ is slightly higher than the inception potential of about 40 mV at $4.5^{o}C$, increasing as the temperature increases. Also, note, that this polarization current changes from outward to incoming as the membrane polarization current crosses the capacitive polarization current. The chloride reversal potential in the axon at rest is reported to be about 15 mV. One should also probably expect a small contribution from polarization channels below the inception potential while in dynamical equilibrium with other players in the AP foot: potassium and sodium permeability, chloride channels current and sodium-potassium pump. However, the present analysis doesn’t provide further insight into the details of the $J_{K}$ current components.
@@
Rosenthal-Bezanilla [43] data covers a range of temperatures from $1^{o}$C to $30^{o}$C yielding the exponential temperature dependence for parameters of interest in our phenomenological theory. Passage of ions through a membrane channel involves interactions with the surface of the pore at microscopic distances. Plots of ionic time rates as Boltzmann’s kinetic rates yield temperature independent Arrhenius activation energies for the passage of ions through the membrane (See Fig. 13). These energies are of the same order as the average of 0.65 eV rate-limiting metabolic biochemical reactions [15]. Traditional Arrhenius processes rely on thermal activation, where temperature plays a direct role in overcoming an energy barrier. Our observed temperature independent Arrhenius rates suggest that thermal fluctuations are not the driving factor in the passage of ions through channels. If the activation energy (typically defined by an Arrhenius relationship) appears constant and unchanging with temperature, it might suggest an alternative mechanism, such as: QuantumQuantum tunnelingtunneling: This would allow ions to ”tunnel” through energy barriers without requiring thermal energy. It can be argued that the tunneling rates are equal to ionic time rates $\mu_{X}$. FixedFixed fieldfield or electromagneticelectromagnetic activationactivation: In the presence of a strong electric field, ion movement might be governed by field strength rather than thermal fluctuations. In these cases, it’s conceivable that $\alpha$, as suggested by our fittings, as a measure of the electromagnetic interaction strength, could be more relevant than thermal energy in setting the ”activation” characteristics for the process.
@@
3. The ions traverse ionic pores interacting with the surface of the pore at microscopic distances. Our fittings of the fraction of open channels with mAvrami equations yield the phenomenological role of the fine-structure constant, the Arrhenius activation energies that are independent of temperature and the role of the temperature dependent ionic time rates. These facts point to the possibility of an explicit quantum tunnelling theory justifying the empirical mAvrami equation.
Presented results will provide a framework for further experimental analysis of excitability role by sodium channels lattice and its thermodynamic phase changing behaviour, the role of quantum mechanics mediating the flow of ions across ion channels and the plasticity of neuronal stable states essential for memory storage and retrieval mediated by continuous phase changes.